OpenStudy (anonymous):

If f(x) = 2x + 3 and g(x) = 4x – 5, what is f(x) ◦ g(x)? 18x + 3 8x – 7 8x – 15 8x + 7

OpenStudy (anonymous):

B

jimthompson5910 (jim_thompson5910):

f(x) = 2x + 3 f((g(x)) = 2g(x) + 3 f((g(x)) = 2(4x – 5) + 3 f((g(x)) = 8x – 10 + 3 f((g(x)) = 8x – 7

OpenStudy (anonymous):

rewrite it as F(G(X)) or 2(4x-5)+3 8x-10+3 =8x-7

OpenStudy (anonymous):

answer is B